# Projected state ensemble of a generic model of many-body quantum chaos

### v1

@@ I Introduction
The problem of thermalization has been a central topic in the study of many-body systems in physics in recent years. In the framework of closed dynamics, the conventional understanding of thermalization is based on the idea that a large many-body system can act as its own thermal bath, leading to relaxation towards thermal equilibrium when any finite subportion is kept in contact with an infinitely large complementary system Cazalilla and Rigol (2010); Polkovnikov et al. (2011); D’Alessio et al. (2016). This paradigm has been extensively studied and its origin has been related to the fundamental properties of the statistics of eigenstates, showing the importance of a probabilistic approach in complex quantum systems, ultimately leading to the use of random matrices Mehta (2004) and the notion of eigenstate thermalisation hypothesis Deutsch (1991); Srednicki (1994); Rigol et al. (2008). The emergence of thermalisation has been verified in several numerical Kim et al. (2014) and even experimental platforms Trotzky et al. (2012); Langen et al. (2013); Geiger et al. (2014); Langen et al. (2015); Neill et al. (2016); Clos et al. (2016); Kaufman et al. (2016). In most cases, the experimental observation of a subsystem is naturally associated to the measurement of its complement, an information that is usually discarded in the standard approach to thermalisation. Thus, for both fundamental and practical motivations, it becomes interesting to explore a thermalization process when the bath is subject to projective measurements and the measurement outcomes are known and stored. This scenario leads to the notion of projected ensemble (PE) Cotler et al. (2023); Choi et al. (2023), where after preparing a state through unitary evolution, one focuses on the ensemble of pure states in the subsystem $A$ conditioned on the outcomes of quantum measurements in the complement $B$ (bath). The PE can be interpreted as a particular unraveling of the reduced density matrix, which is recovered if the information about $B$ is discarded. Thus, going beyond the characterisation of the reduced density matrix, one introduces deep thermalization Ippoliti and Ho (2023); Shrotriya and Ho (2023), whose goal, in broad terms, is to determine the minimal amount of information required to fully characterize this projected ensemble. Arguing on the basis of ergodicity and tipicality of many-body quantum systems, it is reasonable to assume that, for generic chaotic systems and in the absence of specific local conserved quantities, the PE approaches a uniform distribution over the set of pure states in $A$, known as the Haar ensemble. Such a natural conjecture has been confirmed by numerical and experimental evidence in Cotler et al. (2023); Choi et al. (2023). More analytical and even rigorous results provided in Refs. Ho and Choi (2022); Claeys and Lamacraft (2022); Ippoliti and Ho (2023) have further substantiated these claims for classes of strongly chaotic yet tractable models in $1D$, known as dual-unitary circuits (DUC) Bertini et al. (2019a, b); Piroli et al. (2020a). The fact that an ensemble of pure states provides a good approximation of the uniform Haar distribution is an important characteristic known as quantum state design Renes et al. (2004); Ambainis and Emerson (2007). More specifically, consider an ensemble of pure states =$\mathcal{E}=\{p(i),\ket{\Psi(i)}\in\mathcal{H}\}_{i}$, where $p(i)$ is the probability of picking the pure state |$\ket{\Psi(i)}$, labeled by $i$, in the Hilbert space $\mathcal{H}$. Then, $\mathcal{E}$ is dubbed a $k$-design if it provides a good approximation, up to a small $\epsilon$, of the Haar distribution up to $k$ tensor powers of the state itself, i.e.
@@
In this work, we extend the study of deep thermalization to the case of extended random unitary circuits (RUC). RUC are constructed by evolving a spin chain using local gates drawn from some convenient distribution, e.g. the Haar measure over the unitary group, and acting between neighboring spins Fisher et al. (2023). By leveraging the fact that each gate is drawn from the Haar ensemble, the average of many relevant quantities naturally requires an effective pairing of degrees of freedom in spacetime Nahum et al. (2017), eventually leading to a mapping onto a classical statistical mechanics model. The resulting classical models often become tractable at least in the large-dimensional limit of the local Hilbert space, $q\to\infty$. This approach paved the way toward the study of operator spreading, a concept directly related to the butterfly effect and out-of-time order correlators Nahum et al. (2018); von Keyserlingk et al. (2018b); Maldacena and Stanford (2016). Furthermore, RUCs provided a solvable example of the growth of bipartite entanglement in chaotic circuits in the absence of quasiparticles Kim and Huse (2013a). More precisely, entanglement growth can be elegantly and effectively reinterpreted as the cost associated with a membrane (or domain wall) that separates different types of pairing in space-time Nahum et al. (2017). It has been argued that this formulation of entanglement in terms of a membrane is a general feature of chaotic systems, beyond the RUC in which it was actually derived Zhou and Nahum (2020); Chan et al. (2018); Rakovszky et al. (2019) and brought out the role of the Kardar-Parisi-Zhang (KPZ) equation Corwin (2012) in describing its universal fluctuations in generic noisy dynamics Zhou and Nahum (2019); Knap (2018). More recently, RUC have been employed in the study of monitored dynamics, in situations in which unitary evolution is flanked by projective measures. This treatment has made it possible to conjecture the existence of measurement-induced phase transitions, which separate a volume law phase from an area law as the rate of projective measurements increases Skinner et al. (2019); Li et al. (2018, 2019); Cao et al. (2019). In contrast to these cases, in this paper we focus on the aforementioned projected ensemble, where the simultaneous measurements of all sites in $B$ follows the purely unitary dynamics of the whole system, thus revealing a pure state in $A$. We stress that we are interested in the PE exclusively generated by measurements for a fixed realisation of the RUC performing the unitary dynamics. However, to take advantage of the Haar average of the circuit, we focus on the calculation of the frame potential $F^{(k)}_{\mathcal{E}}$, a quantity that measures the $k$-th moment of overlap between two random states in the ensemble $\mathcal{E}$. One can show (see below) that, among all ensembles, the Haar one provides the minimal frame potential, so that $F^{(k)}_{\rm PE}\geq F^{(k)}_{\rm Haar}$. Thus, because of such an exact inequality, the frame potential can be averaged over the realisation of the RUC, leading to an estimate of the distance from a $k$-design of each individual realisation. Remarkably, we found that the calculation of circuit-averaged frame potential leads to a model of statistical mechanics in the presence of domain-wall boundary conditions, similarly to what happens in the calculation of entanglement entropies. In this context, we can once again link the emergence of a $k$-design and deep thermalization to the existence of a membrane: at short times, the membrane starts as a vertical structure with a free end point exiting on the initial factorized state. However, as time progresses, it is less costly for the membrane to be pinned on the boundary, and the value of the frame potential saturates. In particular, for large $q$, the calculation of the frame potential in the projected ensemble can be exactly linked to the moments of the purity of the subregion $A$. Building on this membrane picture in generic models, we explore the saturation regime and estimate the behavior of the design times $t_{k}$. We discover that the logarithmic dependence (LABEL:eq:designtimes), originally found in the toy model of Ippoliti and Ho (2022), is thus quite robust and has its origin in the exponential suppression of the unpinned configurations for the membrane at times $t\gtrsim L_{A}$.
@@ II.1 Frame potential
$\rho^{(k)}_{\mathcal{E}}:=\sum_{i}p(i)\ket{\Psi(i)}\bra{\Psi(i)}^{\otimes k}\;.$ (4)
@@ II.3 Projected state ensemble
The projected ensemble (PE) is based on the following idea: Given a quantum state |$\ket{\Psi}$, belonging to a Hilbert space which can be partitioned into two complementary regions $A$ and $B$, we can measure a complete set of local observables on $B$, and look at the resulting pure state on $A$. More specifically, the ensemble of projected states is given as a set of pairs of probabilities and states
@@ II.4 Replica trick
Since in Eq. (18) the exponent $n$ is simply associated with the norm of each state, in what follows, we will use the nomenclature of spectators for the $n$ replicas, as opposed to the remaining $k$ whose overlap is calculated. In Appendix A, as an instructive example, we provide the evaluation of the projected ensemble for the Haar state ensemble.
@@ III.1 Model
We now consider the projected ensemble starting from a state which is generated by a realisation of the brick-wall random circuit model (BWM). The BWM is a random quantum circuit with the brick-wall geometry, defined as
@@ III.2 Set-up
which can be seen as a classical statistical mechanics model on a strip where the local degrees of freedom are permutations and the Boltzmann’s weights are controlled by the Weingarten’s functions (25) and the scalar products (27). The weights have an intricate structure and can be negative, which complicates the analytical calculation needed for the replica limit $n\to 1-k$. For this reason, we proceed by considering the large-$q$ limit.
@@ IV Mapping between frame potential and purity at large $q$
Figure 3: Diagrammatical illustration of the frame potential $F^{(k,n)}$ where time and space run vertically and horizontally respectively, and the system size of region $A$ is much smaller than the one of region $B$, i.e. $L_{A}\ll L_{B}$. (a) For $t\leq L_{A}/2$, the frame potential is a sum of configurations with $k$ uncorrelated random walks of domain walls. (b) For $t>L_{A}/2$, the frame potential is the $k$ copies of the shortest walks from the boundary of $A$ and $B$, towards the opposite boundary of $A$.
@@
where $\tau_{ij}$ is the transposition permutation exchanging $i$ with $j$.
@@
We see the analogy emerging at large $q$ between the calculation of the frame potential $\overline{F^{(k,n)}_{U(t)}}$ needed for the projected ensemble and $\overline{(F^{(1)}_{U(t)})^{k}}$ which instead leads to the moments of the purity. In particular, a intrinsic property of both computations is that one is left with a domain-wall boundary condition with states in $A$ and $B$ differing by a product of $k$ commuting transpositions. In this case, in leading order in $q\to\infty$, the $k$ domain walls behave as independent random walks; see Fig. 3. Following the derivation in Chan et al. (2018), we write
@@
This result shows that in the leading order at large $q$ the saturation of the frame potentials occurs with a sharp transition for all design times $t_{k}=L_{A}/2$ irrespectively of $k$. A similar behavior has already been observed in dual-unitary circuits Ho and Choi (2022); Ippoliti and Ho (2023); Claeys and Lamacraft (2022) and is thus generally expected in extremely chaotic quantum many-body systems.
@@ V Membrane picture at finite $q$
At large times $v_{2}t\gg L_{A}$, the saturation $F_{\rm BWM}^{(k)}\to F_{\rm H}^{(k)}\sim k!e^{-ks_{\rm eq}L_{A}}$ is the consequence of all $k$ domain walls exiting the left boundary as in Fig. 3b. The decay rates $v^{(k)}$ result from the interaction of the walls of the $k$ domain, and we expect that again the $k$ dependence of $v^{(k)}$ is the result of a fluctuating interface, as described by the KPZ equation. However, as discussed in Zhou and Nahum (2019), the crossover length $\ell_{d}=O(q^{4})$ is rather long even at $q=2$. In our units, the diffusion constant $D=1/2$ and for times $t\lesssim\ell_{d}^{2}$, the $k$ domain walls behave as independent random walks, as the interactions between them are not yet relevant. Consequently, for relatively large values of $L_{A}$ saturation in Eq. (50) occurs well before the effect of interaction plays any role, and we can simply approximate $v^{(k)}\sim k\tilde{v}_{2}$. A quantitative comparison between Eq. (50) and numerical simulations at $q=2$ is shown in Fig. 4.
Figure 4: Numerical simulation of the frame potential $F^{(k)}_{\rm BWM}$ for $q=2$ and $L_{A}=6$ with $k=1,2,3$ and $L_{B}=4,6,\ldots,16$ (black to orange). Comparison with Eq. (50) with $v^{(k)}=k\tilde{v}_{2}$ is shown in red, while the Haar saturation value is plotted in blue. Case $k=1$ is included as a benchmark, since the short-time decrease is exactly given by the average purity (49).
@@ VI Saturation and design times
A finite $q$, not only the velocity $v^{(k)}$ values in Eq. (50) are modified, but also the saturation at $t\approx L_{A}$ is rounded with respect to the sharp step in Eq. (46). As we saw earlier, in calculating the frame potential using the transfer matrix, we can reduce to summing over the configurations of $k$ paths. Moreover, in practice for the accessible values of $L_{A}$, we can neglect the interactions between the paths. Therefore, near the saturation that occurs at times $t=O(L_{A})$, the relevant configurations for each path are reduced to the one that ends at the left or the bottom boundary. For instance, for $k=1$, one has using Zhou and Nahum (2019)
@@
Since the paths can be assumed to be non-interacting, this is easily generalised to $k>1$, leading to
$\left[\Delta^{(k)}\right]^{2}\sim\left[\Delta^{(k)}\right]^{2}_{\rm non-int}=\left(1+\exp[(-\tilde{v}_{2}t+L_{A})s_{\rm eq}]\right)^{k}-1\;.$ (52)

### v2

@@ I Introduction
The problem of thermalization has been a central topic in the study of many-body systems in physics in recent years. In the framework of closed dynamics, the conventional understanding of thermalization is based on the idea that a large many-body system can act as its own thermal bath, leading to relaxation towards thermal equilibrium when any finite subportion is kept in contact with an infinitely large complementary system Cazalilla and Rigol (2010); Polkovnikov et al. (2011); D’Alessio et al. (2016). This paradigm has been extensively studied and its origin has been related to the fundamental properties of the statistics of eigenstates, showing the importance of a probabilistic approach in complex quantum systems, ultimately leading to the use of random matrices Mehta (2004) and the notion of eigenstate thermalisation hypothesis Deutsch (1991); Srednicki (1994); Rigol et al. (2008). The emergence of thermalisation has been verified in several numerical Kim et al. (2014) and even experimental platforms Trotzky et al. (2012); Langen et al. (2013); Geiger et al. (2014); Langen et al. (2015); Neill et al. (2016); Clos et al. (2016); Kaufman et al. (2016). In most cases, the experimental observation of a subsystem is naturally associated to the measurement of its complement, an information that is usually discarded in the standard approach to thermalisation. Thus, for both fundamental and practical motivations, it becomes interesting to explore a thermalization process when the bath is subject to projective measurements and the measurement outcomes are known and stored Anza and Crutchfield (2021). This scenario leads to the notion of projected ensemble (PE) Cotler et al. (2023); Choi et al. (2023), where after preparing a state through unitary evolution, one focuses on the ensemble of pure states in the subsystem $A$ conditioned on the outcomes of quantum measurements in the complement $B$ (bath). The PE can be interpreted as a particular unraveling of the reduced density matrix, which is recovered if the information about $B$ is discarded. Thus, going beyond the characterisation of the reduced density matrix, one introduces deep thermalization Ippoliti and Ho (2023); Shrotriya and Ho (2023), whose goal, in broad terms, is to determine the minimal amount of information required to fully characterize this projected ensemble. Arguing on the basis of ergodicity and typicality of many-body quantum systems, it is reasonable to assume that, for generic chaotic systems and in the absence of specific local conserved quantities, the PE approaches a uniform distribution over the set of pure states in $A$, known as the Haar ensemble. Such a natural conjecture has been confirmed by numerical and experimental evidence in Cotler et al. (2023); Choi et al. (2023). More analytical and even rigorous results provided in Refs. Ho and Choi (2022); Claeys and Lamacraft (2022); Ippoliti and Ho (2023) have further substantiated these claims for classes of strongly chaotic yet tractable models in $1D$, known as dual-unitary circuits (DUC) Bertini et al. (2018, 2019a, 2019b); Piroli et al. (2020a). The fact that an ensemble of pure states provides a good approximation of the uniform Haar distribution is an important characteristic known as quantum state design Renes et al. (2004); Ambainis and Emerson (2007). More specifically, consider an ensemble of pure states =$\mathcal{E}=\{p(i),\ket{\Psi(i)}\in\mathcal{H}\}_{i}$, where $p(i)$ is the probability of picking the pure state |$\ket{\Psi(i)}$, labeled by $i$, in the Hilbert space $\mathcal{H}$. Then, $\mathcal{E}$ is dubbed a $k$-design if it provides a good approximation, up to a small $\epsilon$, of the Haar distribution up to $k$ tensor powers of the state itself, i.e.
@@
In this work, we extend the study of deep thermalization to the case of spatially-extended random unitary circuits (RUC). RUC are constructed by evolving a spin chain using local gates drawn from some convenient distribution, e.g. the Haar measure over the unitary group, and acting between neighboring spins Fisher et al. (2023). By leveraging the fact that each gate is drawn from the Haar ensemble, the average of many relevant quantities naturally requires an effective pairing of degrees of freedom in spacetime Nahum et al. (2017), eventually leading to a mapping onto a classical statistical mechanics model. The resulting classical models often become tractable at least in the large-dimensional limit of the local Hilbert space, $q\to\infty$. This approach paved the way toward understanding spectral statistics Bohigas et al. (1984); Chan et al. (2018a, b, 2019); Garratt and Chalker (2021), which, beyond the late-time emergence of the random matrix theory, displays early-time universal deviation Chan et al. (2018b, 2021); Shivam et al. (2023), and operator spreading, a concept directly related to the butterfly effect and out-of-time order correlators Nahum et al. (2018); von Keyserlingk et al. (2018b); Maldacena and Stanford (2016). Furthermore, RUCs provided a solvable example of the growth of bipartite entanglement in chaotic circuits in the absence of quasiparticles Kim and Huse (2013a). More precisely, entanglement growth can be elegantly and effectively reinterpreted as the cost associated with a membrane (or domain wall) that separates different types of pairing in space-time Nahum et al. (2017). It has been argued that this formulation of entanglement in terms of a membrane is a general feature of chaotic systems, beyond the RUC in which it was actually derived Zhou and Nahum (2020); Chan et al. (2018a); Rakovszky et al. (2019) and brought out the role of the Kardar-Parisi-Zhang (KPZ) equation Corwin (2012) in describing its universal fluctuations in generic noisy dynamics Zhou and Nahum (2019); Knap (2018). More recently, RUC have been employed in the study of monitored dynamics, in situations in which unitary evolution is flanked by projective measures. This treatment has made it possible to conjecture the existence of measurement-induced phase transitions, which separate a volume law phase from an area law as the rate of projective measurements increases Skinner et al. (2019); Li et al. (2018, 2019); Cao et al. (2019). In contrast to these cases, in this paper we focus on the aforementioned projected ensemble, where the simultaneous measurements of all sites in $B$ follows the purely unitary dynamics of the whole system, thus revealing a pure state in $A$. We stress that we are interested in the PE exclusively generated by measurements for a fixed realisation of the RUC performing the unitary dynamics. However, to take advantage of the Haar average of the circuit, we focus on the calculation of the frame potential $F^{(k)}_{\mathcal{E}}$, a quantity that measures the $k$-th moment of overlap between two random states in the ensemble $\mathcal{E}$. One can show (see below) that, among all ensembles, the Haar one provides the minimal frame potential, so that $F^{(k)}_{\rm PE}\geq F^{(k)}_{\rm Haar}$. Thus, because of such an exact inequality, the frame potential can be averaged over the realisation of the RUC, leading to an estimate of the distance from a $k$-design of each individual realisation. Remarkably, we found that the calculation of circuit-averaged frame potential leads to a model of statistical mechanics in the presence of domain-wall boundary conditions, similarly to what happens in the calculation of entanglement entropies. In this context, we can once again link the emergence of a $k$-design and deep thermalization to the existence of a membrane: at short times, the membrane starts as a vertical structure with a free end point exiting on the initial factorized state. However, as time progresses, it is less costly for the membrane to be pinned on the boundary, and the value of the frame potential saturates. In particular, for large $q$, the calculation of the frame potential in the projected ensemble can be exactly linked to the moments of the purity of the subregion $A$. Building on this membrane picture in generic models, we explore the saturation regime and estimate the behavior of the design times $t_{k}$. We discover that the logarithmic dependence, originally found in the toy model of Ippoliti and Ho (2022), is thus quite robust and has its origin in the exponential suppression of the unpinned configurations for the membrane at times $t\gtrsim L_{A}$.
@@ II.1 Frame potential
$\rho^{(k)}_{\mathcal{E}}:=\sum_{i}p(i)\left(\ket{\Psi(i)}\bra{\Psi(i)}\right)^{\otimes k}\;,$ (4)
@@ II.3 Projected state ensemble
The projected ensemble (PE) is based on the following idea: Given a quantum state |$\ket{\Psi}$, belonging to a Hilbert space which can be partitioned into two complementary regions $A$ and $B$, we can measure a complete set of local observables on $B$, and look at the resulting pure state on $A$ (see for example Fig. 1). More specifically, the ensemble of projected states is given as a set of pairs of probabilities and states
@@ II.4 Replica trick
Since in Eq. (18) the exponent $n$ is simply associated with the norm of each state, in what follows, we will use the nomenclature of spectators for the $n$ replicas, as opposed to the remaining $k$ whose overlap is calculated.
@@ III.1 Model
We now consider the projected ensemble starting from a state which is generated by a realisation of the brick-wall random circuit model (BWM), as illustrated in Fig. 1. The BWM is a random quantum circuit with the brick-wall geometry, defined as
@@ III.2 Set-up
which can be seen as a classical statistical mechanics model on a strip where the local degrees of freedom are permutations and the Boltzmann’s weights are controlled by the Weingarten’s functions (25) and the scalar products (27). The weights have an intricate structure and can be negative, which complicates the analytical calculation needed for the replica limit $n\to 1-k$. For this reason, we proceed by considering the large-$q$ limit. Additionally, with the notations introduced, in Appendix A we provide the evaluation of the projected ensemble for the Haar state ensemble as an instructive example.
@@ IV Mapping between frame potential and purity at large $q$
Figure 3: Illustration of the frame potential $F^{(k,n)}$ where time and space run vertically and horizontally respectively, and the system size of region $A$ is much smaller than the one of region $B$, i.e. $L_{A}\ll L_{B}$. (a) For $t\leq L_{A}/2$, the frame potential is a sum of configurations with $k$ uncorrelated random walks of domain walls. (b) For $t>L_{A}/2$, the frame potential is the $k$ copies of the shortest walks from the boundary of $A$ and $B$, towards the opposite boundary of $A$. Note that, at large local Hilbert space size $q$, $\mu_{A}$ is a fixed permutation in $S_{2n+2k}$, while $\mu_{B}(\alpha_{k})$ can assume $k!$ values according to some permutation $\alpha_{k}\in S_{k}$, see Eq. (40).
@@
where $\tau_{ij}$ is the transposition permutation exchanging $i$ with $j$. Therefore, to highlight the fact that $\mu_{B}$ depends on $\alpha_{k}\in S_{k}$, we write $\mu_{B}(\alpha_{k})$.
@@
We see the analogy emerging at large $q$ between the calculation of the frame potential $\overline{F^{(k,n)}_{U(t)}}$ needed for the projected ensemble and $\overline{(F^{(1)}_{U(t)})^{k}}$ which instead leads to the moments of the purity. In particular, an intrinsic property of both computations is that one is left with a domain-wall boundary condition with states in $A$ and $B$ differing by a product of $k$ commuting transpositions. In this case, in leading order in $q\to\infty$, since the permutations at the top boundary of frame potential factorize in Fig. 2, the $k$ domain walls behave as independent random walks, see Fig. 3. Following the derivation in Chan et al. (2018a), we write
@@
This result shows that in the leading order at large $q$ the saturation of the frame potentials occurs with a sharp transition for all design times $t_{k}=L_{A}/2$ irrespectively of $k$. As we will see below, the sharp transition at $t=O(L_{A})$ will smoothen out at finite values of $q$ due to comparable contributions from configurations where walks exit at the bottom or left boundaries (see Fig. 3). A similar behavior to Eq. (46) has been observed in dual-unitary circuits Ho and Choi (2022); Ippoliti and Ho (2023); Claeys and Lamacraft (2022) and is thus generally expected in extremely chaotic quantum many-body systems.
@@ V Membrane picture at finite $q$
At large times $v_{2}t\gg L_{A}$, the saturation $F_{\rm BWM}^{(k)}\to F_{\rm H}^{(k)}\sim k!e^{-ks_{\rm eq}L_{A}}$ is the consequence of all $k$ domain walls exiting the left boundary as in Fig. 3b. The decay rates $v^{(k)}$ result from the interaction of the walls of the $k$ domain, and we expect that again the $k$ dependence of $v^{(k)}$ is the result of a fluctuating interface, as described by the KPZ equation. However, as discussed in Zhou and Nahum (2019), the crossover length $\ell_{d}=O(q^{4})$ is rather long even at $q=2$. In our units, the diffusion constant $D=1/2$ and for times $t\lesssim\ell_{d}^{2}$, the $k$ domain walls behave as independent random walks, as the interactions between them are not yet relevant. Consequently, for relatively large values of $L_{A}$, saturation in Eq. (50) occurs well before the effect of interaction plays any role, and the sum over paths involved in the frame potential factorises as $k$ independent paths, thus implying that we can approximate $v^{(k)}\sim k\tilde{v}_{2}$. A quantitative comparison between Eq. (50) and numerical simulations at $q=2$ is shown in Fig. 4. At the accessible system sizes, it is difficult to accurately estimate the slope $v^{(k)}$ and its dependence on $k$, but we observe that $v^{(k)}<k\tilde{v}_{2}$ and the difference increases with $k$. This is consistent with the effect of attractive interactions between replicas Zhou and Nahum (2019). In other words, $F^{(k)}_{BWM}$ is larger at finite $q$ because multiple domain walls correlate by following the same favourable path. This gives an additional hint about the emergence of KPZ physics also for this quantity.
Figure 4: Numerical simulation of the frame potential $F^{(k)}_{\rm BWM}$ for $q=2$ and $L_{A}=6$ with $k=1,2,3$ and $L_{B}=4,6,\ldots,16$ (black to orange). Comparison with the pre-saturation regime of Eq. (50) is shown in red, while the Haar saturation value is plotted in blue. Case $k=1$ is included as a benchmark, since the short-time decrease is exactly given by the average purity (49). Note that the non-interacting approximation overestimates the slope and the error increases with $k$: a numerical fit gives the estimates $v^{(k)}=k\tilde{v}_{2}-\delta v^{(k)}$, with $\delta v^{(2)}\sim 0.7$ and $\delta v^{(3)}\sim 0.4$.
@@ VI Saturation and design times
A finite $q$, not only the speed $v^{(k)}$ values in Eq. (50) are modified, but also the saturation at $t\approx L_{A}$ is rounded with respect to the sharp step in Eq. (46). As we saw earlier, in calculating the frame potential using the transfer matrix, we can reduce to summing over the configurations of $k$ paths. Moreover, in practice for the accessible values of $L_{A}$, we can neglect the interactions between the paths. Therefore, near the saturation that occurs at times $t=O(L_{A})$, the relevant configurations for each path are reduced to the one that ends at the left or the bottom boundary. For instance, for $k=1$, one has using Zhou and Nahum (2019)
@@
This is easily generalised to $k>1$. Indeed, since the paths can be assumed to be non-interacting, the $k$ paths can independently choose whether to exit from the left or the bottom, i.e. $F^{(k)}_{\rm BWM}\sim[F^{(1)}_{\rm BWM}]^{k}$ for $k>1$. Consequently,
$\left[\Delta^{(k)}\right]^{2}\sim\left(1+\exp[(-\tilde{v}_{2}t+L_{A})s_{\rm eq}]\right)^{k}-1=:\left[\Delta^{(k)}\right]^{2}_{\rm non-int}\;.$ (52)
Note that Eq. (52) describes the regime $t=O(L_{A})$ where the contributions of the configurations with paths exiting the left or the bottom are comparable. In the regimes where $t\ll L_{A}$ and $t\gg L_{A}$, Eq. (52) is consistent with Eq. (50).